OpenStudy (anonymous):

Which of the following is equivalent to |4x + 5| = 88?

OpenStudy (anonymous):

\[ 4 x+5 = \pm 88 \]

OpenStudy (anonymous):

so -93/4=x or 83/4=x

OpenStudy (anonymous):

4x+5=88 4x+5=-88

OpenStudy (anonymous):

solve for both equations.